Subject: APO 3.5m users committee minutes, April 2014

From: Suzanne Hawley

Submitted: Wed, 23 Apr 2014 14:41:23 -0700 (PDT)

Message number: 1331 (previous: 1330, next: 1332 up: Index)

 		APO 3.5-m Users Committee Phonecon, 4/14/14

Attending: Suzanne Hawley, Mark Klaene, Scott Anderson, Rachel Kuzio de 
Naray, Bill Ketzeback, Bruce Gillespie, Jeff Bary, Rene Walterbos, Russet 
McMillan, John Bally, Michael Strauss, Mary Beth Kaiser, Joanne Hughes 
Clark, Remy Indebetouw, Jon Holtzmann

Absent: Al Harper, John Wisniewski


User feedback and comments from institutional representatives:

o Washington - Scott had nothing to report.

o Chicago - No report.

o New Mexico State - Jon had nothing to report.

o Princeton - Michael said Princeton is happy.

o Colorado - John said thanks for the recent excellent student field trip, 
and they got some great data.

o Virginia - Remy had nothing to report.

o Johns Hopkins - Mary Beth had nothing to report.

o NAPG - Jeff said that there was a group of students that had a 
successful APO visit with Beth Willman from Haverford.

o Georgia State - Rachel had nothing to report.

o Seattle Univ. - Joanne had nothing to report.

o Oklahoma - no report.


Discussion of telescope/instruments report:

Mark mentioned the following while discussing the report.

We had a nearby (about 4 or 5 miles north and west of Sunspot) wildfire 
start (most probably human-caused) this past Thursday that grew to about 
100 acres. The fire was contained and is now essentially extinguished. 
There were minor issues with smoke that were monitored closely.  On the 
telescope, we have been refining the new guiding code and it is now in 
regular use, including with Agile.  Thanks go to all the users who helped 
us wring out the new code during their science time. It would have taken 
considerably more engineering time if that had been the only time we could 
have used for the tests.

Instruments have not had any new problems this month. The SPIcam dark 
spikes are still with us, so we plan to service the ion pump during the 
period late April through early May .  Thus SPIcam will be down for 2-3 
weeks. It is not scheduled to be on the telescope during that time, but if 
this poses a problem for any users, please notify Bill Ketzeback right 

The GIFS software (TUI, ICC, hub mods, etc.) is about 90% completed. There 
will be a GIFS training run 27 and 28 May. Anyone who is interested in 
participating should contact Carol Grady and Bill Ketzeback.

The CCD for the ARCTIC imager was chosen to be an STA detector, see the 
ARCTIC web page for details.  Joe Huehnerhoff is finalizing the optical 
design for the instrument, and also carrying out lifecycle tests on the 
new shutter.

We are planning two summer shutdown periods this year, the first is July 
7-24 for normal maintenance, and the second is August 8-17 to install and 
test the TCC upgrade.

Mark Klaene's detailed report for the reporting period follows:


 	      3.5-m Telescope and Instruments Highlights
 	                03/06/14 through 04/09/14

0) Overview

Generally another productive period with variable cloudiness, little 
precipitation.  Current forest fire danger is just moderate. 
Preparations are underway for 20-yr anniversary event.  Housing continues 
to be tight through May due to more training sessions and SDSS-IV work. 
We had a successful class trip from CU this period, thanks to John Bally. 
This class was much larger than we recommend but it worked out well.

1) Telescope

Telescope has been working very well.  Refinements to the new guider code 
are essentially complete.  No further development is planned, however, we 
will address issues that arise.  Fritz put a lot of effort into this and 
it shows.  The new guider code addresses many problems reports over the 
years and is now maintainable.  The new TCC development is also 
progressing nicely with some initial testing on the telescope during the 

2) Instruments

DIS has been operating normally.

Agile is operational with no problems.

TripleSpec had no operational problems this period.

SPIcam variable dark current has been isolated to the current ion pump 
type and controller.  Currently we see a spike that may affect dark 
current about 1/day and if it occurs during observations we will let the 
observer know to check their data.  We plan to take the instrument offline 
in late April/early May to address this issue.

Echelle is operational with no problems.

NIC-FPS continues to operate in shared-risk observing mode due to 
occasional image corruption issue.  NFS less than 9 is operational.

GIFS is operational.  The small LN2 hold time designed in by GFSC 
continues to be a problem.  The TUI interface is essentially ready and 
undergoing final user testing.


The instrument development is proceeding on schedule.  The CCD has been 
ordered.  Final testing of the prototype shutter is ongoing and should be 
complete in May.  The optics is converging toward a final f/8 design, with 
plans to begin ordering glass in the summer.

4) Summer shutdown

There will be 2 shutdowns scheduled this summer.  The first is from July 
7th to July 24th for M2 and M3 mirror re-coating, mechanical maintenance, 
electronic system upgrades and the majority of instrument servicing.  A 
second shutdown will be from August 8th to August 17th for the new TCC 
upgrade.  We expect the telescope to be operational for much of the time 
during the second shutdown as is necessary to test the new TCC.


Recent class trips - Klaene:

We recently hosted 3 group visits at APO, which taxed our housing capacity 
but in the end worked out fine.  John Bally noted that staging his 
activities from the conference area at the Sunspot visitor center was a 
boon for his large group of students.


ARCSAT shared-risk observing - Hawley, Ketzeback:

Suzanne thanked Joe Huehnerhoff and Bill Ketzeback, along with other APO 
staff including Ed Leon, Ben Harris, Conor Sayres, and Russell Owen, for 
their hard work in bringing the ARCSAT upgrade project to successful 
completion. More than ten different users have now tested the telescope, 
including a group of students from U. Oklahoma who used it while on-site 
for a training trip.  We have gotten good feedback on the operation of the 
ACP software interface, and are still working on a few issues to make it 
fully compatible with our remote observing operations.

We are now at a stage where we can begin general ARCSAT observing in a 
shared-risk mode. Suzanne noted that PhD scientists and graduate students 
at all ARC partners and leasing institutions are welcome to use ARCSAT 
remotely, with no required trip to the site for training.  Undergraduate 
students are also welcome if a PhD/graduate student is present with them 
during observing.  A Call for Proposals for ARCSAT will go out soon for 
the period May 5 - Oct 5, 2014.  The telescope will be scheduled in one 
week blocks (Monday-Sunday), and there will be no charge during this time. 
We will appreciate helpful feedback to improve operations during this 
shared-risk period. See the ARCSAT webpage (linked off the main APO home 
page) for more information, including the users manual, the proposal form 
and the call for proposals.  Proposals for the shared risk period should 
be sent by email (as text files only!) to Suzanne and Russet by May 1, 

Jon Holtzman mentioned that there are limited opportunities for robotic 
programs to be conducted on the NMSU 1-meter telescope; contact Jon if 


WIYN trades update - Hawley:

Suzanne received a couple of proposals for WIYN time trades for Q3/4, 
comprising a few nights of WIYN observing time.  Suzanne said that John 
Salzer, WIYN Board President, felt that the 2014B WIYN observing schedule 
could support these requests, and that Wisconsin and Indiana users are 
looking into using the ARC telescope on trade.  If any users are still 
interested in using WIYN time during the 2014B semester, by trading ARC 
time in Q3/Q4, they should contact Suzanne as soon as possible.


Adler NEO program - Hawley, Walterbos:

Rene was pleased to announce that Adler planetarium has been awarded a 
2-year NASA grant to use the 3.5-m to observe NEOs photometrically and 
spectroscopically. Adler will lease the 17% share of the telescope that is 
being vacated by Chicago at the end of 2014.  This will affect how much 
time is available for new leasing partners in 2015-2016.  Current ARC 
partners and leasing partners with signed 3 year leases will not be 

Leasing and new partnership arrangements for ARC are being discussed by 
telecon every other week by the New Members Committee comprised of Rene 
Walterbos, Suzanne Hawley, Ron Irving, Mike Evans, Bruce Gillespie, and 
Mike Skrutskie.  If users have questions or input to this process, they 
should contact any of these people.


20th anniversary celebration - Hawley, Walterbos:

The 3.5m 20th anniversary (and ARC 30th anniversary) fest will be held at 
APO on 12-13 May.  There will be a reception and star party, a program of 
historical and scientific talks, lunch and a tour on-site and dinner at 
the Lodge.  We are expecting about 75 people to attend the festivities.



[open from previous meeting]:

        ===> ACTION:  Advertise ARCSAT shared-risk observing starting in May
              Status:  Discussed above, closed.

        ===> ACTION:  Send expressions of interest for WIYN time trades to 
Suzanne by April 1
              Status:  Discussed above, closed.

[new from this meeting]:



Next meeting:  The next Users Committee phonecon will be on Monday 19 May 
at 8:30 AM Pacific Time.  The agenda and other materials will be sent to 
the committee members during the preceding week.

APO APO APO APO APO  Apache Point Observatory 3.5m  APO APO APO
APO  This is message 1331 in the apo35-general archive. You can find
APO  the archive on
APO  To join/leave the list, send mail to
APO  To post a message, mail it to